Completely Nuts Peanut Butter wanted to analyze their results for their most recent month of sales in the Southwestern US. Their plan and actual results were as follows:

Measure March
Plan
March
Actual
Demand 25,000,000 26,100,000
Share 14.3% 14.5%
Marketing Exp. 300,000 $288,000
Price $3.50 $3.18
Cost $1.70 $1.50

What is the NMC variance (in 000s) for Completely Nuts?
Answer #1

Answer:

  • NMC = Net marketing contribution = (Volume * Margin) - MArketing expenses
  • Where , Volume = Demand * Share
  • Margin = Price - Cost

March Plan :

  • Margin = Price - Cost = 3.5 - 1.7 = $1.80
  • Volume = Demand *Share = 25,000,000 *0.143 = 3,575,000
  • Marketing Exp = 300,000
  • NMC Planned = (3,575,000*1.80) - 300,000 = $6,135,000

March Actual:

  • Margin = Price - Cost = 3.18 - 1.5 = $1.68
  • Volume = Demand *Share = 26,100,000 *0.145 = 3,784,500
  • Marketing Exp = 288,000
  • NMC Actual = (3,784,500*1.68) - 288,000 = $6,135,000 = $ 6,069,960

NMC VAriance = NMC Actual - NMC Planned = 6,069,960- 6,135,000​​​​ = ​​​- $ 65,040 = $65.04 K (Answer)